# Physics of droplet regulation in biological cells

### v2

@@ Abstract
  • Version of draft: February 27, 2025
@@ 1 Introduction
Biological cells contain droplets, known as biomolecular condensates, which partition the cells’ interior without using lipid membranes [1]. Besides simply structuring cells, these droplets also provide distinct chemical environments [2], affect reactions [3], sequester molecules [4], exert forces [5], and can induce phenotypic switching [6]. Droplets are implicated in homeostasis of healthy cells (e.g., in development [7], gene regulation [8, 9], protein quality control [10], and signaling [11]), but also in diseased states [12, 13] (e.g., neurodegenerative diseases [14], and cancer [15]). It is thus not surprising that droplets appear everywhere in cells, from the nucleus [16], to the cytosol [1], to membranes [11], and in all forms of life, encompassing bacteria [17], plants [18], and animals [1]. While it is clear that droplets are crucial for cells, it is less clear how they form and how cells control them.
@@
This review summarises recent work on intracellular phase separation, focusing on the physical processes that allow cells to regulate droplets, i.e., processes that determine when and where droplets form, how large they get, and when they dissolve. We focus less on how droplets affect cellular processes in vivo (reviewed in [10, 26, 8, 27, 28, 29, 30, 22, 31, 32, 9, 33, 34, 7, 35, 3, 18, 11, 36, 37, 15, 38, 12, 39, 40, 5, 41, 14, 42, 43, 44, 45, 46, 47, 48, 49, 24, 50, 51, 52, 16, 17, 53, 54, 55, 56, 13, 57, 1, 58]) or in related in vitro studies (reviewed in [26, 8, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 25, 70, 71, 72, 73, 74, 75, 76, 77]). We also neglect external influences, such as electric and magnetic fields, or gravity, which is only relevant in extremely large cells [78], and do not discuss the detailed (bio-)chemistry that leads to phase separation (see [79, 80, 22, 81, 82] for details). Instead, we assume that such interactions can be captured by a suitable coarse-grained free energy, so that we can describe droplets using continuous concentration fields. Our approach is similar to that taken in other reviews of the theory of phase separation in cells [83, 84, 85, 86, 81, 87, 88, 89, 82, 23, 90], but with a particular focus on droplet regulation through the principles of soft matter physics [91, 92, 93].
@@ 2.1.1  Thermodynamics of liquids
We first consider an isothermal, homogeneous system of volume $V_{\mathrm{sys}}$ filled with a liquid comprised of a solvent and ${N_{\mathrm{c}}}$ additional components. For simplicity, we focus here on incompressible systems, where the molecular volumes $\nu_{i}$ of each component are constant. The generalisation to compressible systems is discussed in appendix A. The composition of the homogeneous system is fully described by the particle counts $N_{i}$ for $i=1,\ldots,{N_{\mathrm{c}}}$ and all remaining space is filled with solvent. Since the system is isothermal and has constant volume, the relevant thermodynamic potential is the Helmholtz free energy $F=V_{\mathrm{sys}}f(\phi_{1},\ldots,\phi_{N_{\mathrm{c}}})$, where $f$ is the free energy density which is a function of the ${N_{\mathrm{c}}}$ volume fractions $\phi_{i}=\nu_{i}N_{i}/V_{\mathrm{sys}}$. We consider here the Flory–Huggins free energy [94, 95],
@@
where the first term accounts for the mixing entropy with $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$ due to incompressibility; see appendix A. We also use the molecular volume $\nu_{0}$ of the solvent to define non-dimensional molecular sizes $\Omega_{i}=\nu_{i}/\nu_{0}$. The last two terms in equation (2.1) specify enthalpic interactions using non-dimensional quantities $w_{i}$ and $\chi_{ij}$, whose particular forms can be determined from detailed theories [81]. Here, $w_{i}$ captures the difference of internal energy between species $i$ and the solvent, but it can also account for external potentials. Such potentials can in principle depend on space, due to, for example, gravity or gradients in a regulatory species [23], but we do not discuss this aspect further in this review. In contrast, the symmetric Flory parameters $\chi_{ij}=\chi_{ji}$ capture interactions: positive values raise the free energy and thus relate to repulsion between species $i$ and $j$, while negative values imply effective attraction. Since the interactions $\chi_{ij}$ are measured relative to solvent molecules, the diagonal terms need not vanish; see appendix A.
@@ 2.1.2  Kinetics of heterogeneous systems
To describe droplet dynamics, we must generalise the theory presented in the previous section to heterogeneous systems. To do this, we assume local equilibrium, neglecting the heat fluctuations which can manifest on long length and time scales in these typically active systems [98]. We can thus assume an isothermal system and define fields for the thermodynamic quantities introduced above. In particular, the system state is now specified by volume fraction fields $\Phi(\boldsymbol{r})=\bigl{(}\phi_{1}(\boldsymbol{r}),\ldots,\phi_{{N_{\mathrm{c}}}}(\boldsymbol{r})\bigr{)}$ and its free energy is given by the functional [99, 100, 101]
@@
where the first term accounts for the local free energy described by equation (2.1), and the second term results from a gradient expansion [99]. The symmetric matrix $\kappa_{ij}$ quantifies how strongly compositional gradients are penalised and generally depends on composition [102, 103, 104]. In the simple case where $\kappa_{ij}$ does not depend on composition, the exchange chemical potentials, $\bar{\mu}_{i}=\nu_{0}\Omega_{i}\delta F_{\mathrm{bulk}}/\delta\phi_{i}$, read
@@
We now derive kinetic equations based on linear non-equilibrium thermodynamics [96], starting from the continuity equation, $\partial_{t}\phi_{i}+\nabla\cdot\boldsymbol{J}_{i}=s_{i}$, where $\boldsymbol{J}_{i}$ denotes spatial fluxes and the source term $s_{i}(\boldsymbol{r})$ accounts for chemical reactions as described in section 5. To separate fluxes due to hydrodynamic flows (associated with momentum transport) from diffusive fluxes driven by gradients in chemical potentials, we split $\boldsymbol{J}_{i}$ into a centre-of-mass velocity field $\boldsymbol{v}=\sum_{i=0}^{N_{\mathrm{c}}}\boldsymbol{J}_{i}$ (assuming equal mass density for all components) and relative fluxes $\boldsymbol{j}_{i}=\phi_{i}(\boldsymbol{v}_{i}-\boldsymbol{v})$, where $\boldsymbol{v}_{i}=\boldsymbol{J}_{i}/\phi_{i}$ is the velocity of component $i$ in the lab frame. Hence,
@@
$\partial_{t}\boldsymbol{v}+\rho\boldsymbol{v}\!\cdot\!\nabla\boldsymbol{v}=\nabla\boldsymbol{\sigma}\;,$ (2.9)
@@
The coupled partial differential equations (2.14) for the fields $\{\phi_{i}\}$ typically permit stationary homogeneous solutions. For instance, without chemical reactions ($s_{i}=0$), each homogeneous state $\phi_{i}(\boldsymbol{r})=\bar{\phi}_{i}$ is stationary because of material conservation. Such a state is stable if the associated Hessian
$H_{ij}$ $=-\nu_{0}\Omega_{i}^{2}\boldsymbol{q}^{2}\sum_{l=1}^{N_{\mathrm{c}}}\Lambda_{il}\left[\frac{\partial^{2}f}{\partial\phi_{l}\phi_{j}}+\kappa_{lj}\boldsymbol{q}^{2}\right]+\frac{\partial s_{i}}{\partial\phi_{j}}$ (2.15)
has only positive eigenvalues for all wave vectors $\boldsymbol{q}$. In contrast, the homogeneous state is unstable if $H_{ij}$ has positive eigenvalues and the largest eigenvalue will typically determine the dominant length scale $2\pi/|\boldsymbol{q}|$ from the associated wavevector $\boldsymbol{q}$. However, since equations (2.14) are nonlinear, such a stability analysis only provides information about the initial dynamics.
@@ 2.1.3  Behaviour of binary liquids
Figure 2.1: Phase behaviour of a binary liquid. (A) Free energy density $f$ given by equation (2.16) as a function of the fraction $\phi$ (blue line) for $\Omega=1$, $\chi=2.5$, and $w=0.2$. The common-tangent construction (black lines) solves equations (2.30) and thus determines the equilibrium fractions $\phi_{\mathrm{out}}$ and $\phi_{\mathrm{in}}$ (black disks) for two values of the Laplace pressure $P_{\gamma}$. The inflection points (orange disks) enclose the spinodal region where $f^{\prime\prime}(\phi)<0$. (B) Exchange chemical potential $\bar{\mu}$ given by equation (2.17) as a function of $\phi$ corresponding to panel A. The blue regions of equal area indicating a Maxwell construction, equivalent to the common-tangent construction. (C) Osmotic pressure $\Pi=\nu^{-1}\phi\bar{\mu}-f$ as a function of $\phi$ corresponding to panel A. (D) $\Pi$ as a function of $\bar{\mu}$ corresponding to panel A. (E, F) Phase diagrams corresponding to panel A highlighting the equilibrium volume fractions (black binodal line) and the spinodal region (within the orange spinodal line) for a symmetric mixture ($\Omega=1$, panel E) and for large solutes ($\Omega=10$, panel F). The critical point is marked by a star and grey areas denotes the region where bi-continuous structures have lower energy than droplets in two dimensions (obtained by comparing interfacial energies of disks and stripes for equal length scales). (G, H) Grand-canonical phase diagram corresponding to panels E and F. The binodals (black lines) mark the first-order transition between dilute and dense phases. Both phases are linearly stable in the spinodal region (between orange dotted lines).
@@
and determines the phase behaviour of the liquid; see figure 2.1. In particular, a homogeneous system with volume fraction $\phi$ is stable with respect to small perturbations when the Hessian given in equation (2.15) is negative, i.e., if ${f^{\prime\prime}(\phi)\propto\bar{\mu}^{\prime}(\phi)>0}$. This occurs when the interaction parameter $\chi$ is smaller than
@@
Homogeneous systems are stable against small fluctuations outside the spinodal region, but they are not necessarily the only stable states or even the state with the lowest free energy. To see whether heterogeneous states (e.g., droplets) are able to form, we next consider the equilibrium conditions (2.5). These conditions can be solved graphically using the common–tangent construction (figure 2.1A), or equivalently the equal–area Maxwell construction (figure 2.1B). Alternatively, the equilibrium point can be read off the $\Pi(\bar{\mu})$ plot shown in figure 2.1D, and the corresponding volume fractions then identified via figures 2.1B–C. For the chemical potential of the binary system given by equation (2.17), these graphical constructions are possible, and the solutions unique, if $\chi>\chi_{*}$. The two fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ correspond to the equilibrium fractions of the dilute and dense phase, respectively outside and inside droplets. Note that $\phi^{(0)}_{\mathrm{out}}$ is also known as the saturation concentration since droplets can only form when it is exceeded. The common–tangent construction also illustrates that the phase-separated state generally has lower free energy than the corresponding homogeneous state with equal average composition $\bar{\phi}$. The two states only have equal free energy in the limit where one phase of the phase-separated state vanishes ($\bar{\phi}=\phi^{(0)}_{\mathrm{out}}$ or $\bar{\phi}=\phi^{(0)}_{\mathrm{in}}$). Taken together, the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ as a function of $\chi$ form the binodal line. Since the phase-separated state does not exist outside the binodal line, this line also bounds the region where the homogeneous phase is globally stable; see figure 2.1. The binodal line can be obtained analytically for the symmetric case,
@@
for $\Omega=1$. In the limit of strong interactions ($\chi\gg 1$), this predicts $\phi^{(0)}_{\mathrm{out}}\approx e^{-\chi}$, which reveals that even a minute increase of the interaction energy by $2k_{\mathrm{B}}T$ (so $\chi$ is increased by $2$) leads to an almost $10$-fold decrease in the saturation concentration $\phi^{(0)}_{\mathrm{out}}$. In contrast, the spinodal branch scales as $\phi^{\mathrm{spin}}_{\mathrm{out}}\approx 1/(2\Omega\chi)$, implying that a linear stability analysis of the homogeneous state is inadequate to study phase separation in detail.
The entire phase diagram can be separated into three distinct regions. For low $\chi$ or dilute solutions ($\phi<\phi^{(0)}_{\mathrm{out}}$ or $\phi>\phi^{(0)}_{\mathrm{in}}$) only homogeneous states persist. Conversely, only the phase-separated state exists in the spinodal region when interactions are sufficiently strong ($\chi>\chi_{\mathrm{spin}}$). Between these two regions, homogeneous and phase-separated states are both locally stable, but the phase-separated state has lower energy; nucleation events thus become possible when thermal fluctuations are included.
@@
which implies an extra term $-\kappa\nabla^{2}\phi$ in the exchange chemical potential $\bar{\mu}$ given in equation (2.17). The kinetics of the binary liquid follow from equation (2.14),

### v3

@@ Abstract
  • Version of draft: September 12, 2025
@@ 1 Introduction
Biological cells contain droplets, known as biomolecular condensates, which partition the cells’ interior without using lipid membranes [1]. Besides simply structuring cells, these droplets also provide distinct chemical environments [2], affect reactions [3, 4], sequester molecules [5], exert forces [6], and can induce phenotypic switching [7]. Droplets are implicated in homeostasis of healthy cells (e.g., in development [8], gene regulation [9, 10], protein quality control [11], and signaling [12]), but also in diseased states [13, 14] (e.g., neurodegenerative diseases [15], and cancer [16]). It is thus not surprising that droplets appear everywhere in cells, from the nucleus [17], to the cytosol [1], to membranes [12], and in all forms of life, encompassing bacteria [18], plants [19], and animals [1]. While it is clear that droplets are crucial for cells, it is less clear how they form and how cells control them.
@@
This review summarises recent work on intracellular phase separation, focusing on the physical processes that allow cells to regulate droplets, i.e., processes that determine when and where droplets form, how large they get, and when they dissolve. We focus less on how droplets affect cellular processes in vivo (reviewed in [27, 11, 9, 28, 29, 30, 31, 23, 32, 33, 10, 34, 35, 8, 36, 4, 19, 12, 37, 38, 16, 39, 13, 40, 41, 6, 42, 15, 43, 44, 45, 46, 47, 48, 49, 50, 25, 51, 52, 53, 17, 18, 54, 55, 56, 57, 14, 58, 1, 59, 60, 61, 62]) or in related in vitro studies (reviewed in [61, 62, 9, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 26, 74, 75, 76, 77, 78, 79, 80, 81]). We also neglect external influences, such as electric and magnetic fields, or gravity, which is only relevant in extremely large cells [82], and do not discuss the detailed (bio-)chemistry that leads to phase separation (see [83, 84, 23, 85, 86] for details). Instead, we assume that such interactions can be captured by a suitable coarse-grained free energy [87], so that we can describe droplets using continuous concentration fields. Our approach is similar to that taken in other reviews of the theory of phase separation in cells [88, 89, 90, 91, 85, 92, 93, 94, 86, 24, 95], but with a particular focus on droplet regulation. In essence, we ask what physical processes (mostly inspired by soft matter physics [96, 97, 98]) can cells use to control droplets.
@@ 2.1.1 Thermodynamics of liquids
We first consider an isothermal, homogeneous system of volume $V_{\mathrm{sys}}$ filled with a liquid comprised of a solvent and $N_{\mathrm{c}}$} additional components. For simplicity, we focus on incompressible systems, where the molecular volumes $\nu_{i}$ of each component are constant. The generalisation to compressible systems is discussed in appendix A. The composition of the homogeneous system is fully described by particle counts $N_{i}$ for $i=1,\ldots,{N_{\mathrm{c}}}$ and all remaining space is filled with solvent ($i=0$). Since the system is isothermal and has constant volume, the relevant thermodynamic potential is the Helmholtz free energy $F=V_{\mathrm{sys}}f(\phi_{1},\ldots,\phi_{N_{\mathrm{c}}})$, where $f$ is a free energy density, which is a function of the $N_{\mathrm{c}}$} volume fractions $\phi_{i}=\nu_{i}N_{i}/V_{\mathrm{sys}}$. We consider a generic form that combines translational entropies with pairwise interactions between components (known as the Flory–Huggins free energy [99, 100]),
@@
where the first term accounts for the mixing entropy with $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$ due to incompressibility; see appendix A. We also use the molecular volume $\nu_{0}$ of the solvent to define non-dimensional molecular sizes $\Omega_{i}=\nu_{i}/\nu_{0}$. The last two terms in equation (2.1) specify enthalpic interactions using non-dimensional quantities $w_{i}$ and $\chi_{ij}$, which include solvent interactions (see equation (A.11)) and whose particular values can be determined from detailed theories [85]. Here, $w_{i}$ captures the difference of internal energy between species $i$ and the solvent, but it can also account for external potentials. Such potentials can in principle depend on space, due to, for example, gravity or gradients in a regulatory species [24], but we do not discuss this aspect further in this review. In contrast, the symmetric Flory parameters $\chi_{ij}=\chi_{ji}$ capture interactions: positive values raise the free energy and thus relate to repulsion between species $i$ and $j$, while negative values imply effective attraction. Since the interactions $\chi_{ij}$ are measured relative to solvent molecules, the diagonal terms need not vanish; see appendix A. The Flory–Huggins free energy given by equation (2.1) only describes the basic behavior of phase separating mixtures, and we discuss more complex interactions in section 3.1.
@@ 2.1.2 Kinetics of heterogeneous systems
To describe droplet dynamics, we must generalise the theory presented in the previous section to heterogeneous systems. To do this, we assume local equilibrium, neglecting the heat fluctuations which can manifest on long length and time scales in typically active biological systems [103]. We can thus assume an isothermal system and define fields for the thermodynamic quantities introduced above. In particular, the system state is now specified by volume fraction fields $\Phi(\boldsymbol{r})=\bigl{(}\phi_{1}(\boldsymbol{r}),\ldots,\phi_{{N_{\mathrm{c}}}}(\boldsymbol{r})\bigr{)}$ and its free energy is given by the functional [104, 105, 106]
@@
without interactions with the system’s boundary, which we will include in equation (4.1). Here, the first term accounts for the local free energy described by equation (2.1), and the second term results from a gradient expansion [104]. The symmetric matrix $\kappa_{ij}$ quantifies how strongly compositional gradients are penalised and generally depends on composition [107, 108, 109]. In the simple case where $\kappa_{ij}$ does not depend on composition, the exchange chemical potentials, $\bar{\mu}_{i}=\nu_{0}\Omega_{i}\delta F_{\mathrm{bulk}}/\delta\phi_{i}$, read
@@
We now derive kinetic equations based on linear non-equilibrium thermodynamics [110, 101], starting from the continuity equation, $\partial_{t}\phi_{i}+\nabla\cdot\boldsymbol{J}_{i}=s_{i}$, where $\boldsymbol{J}_{i}$ denotes spatial fluxes and the source term $s_{i}(\boldsymbol{r})$ accounts for chemical reactions as described in section 5. To separate fluxes due to hydrodynamic flows (associated with momentum transport) from diffusive fluxes driven by gradients in chemical potentials, we split $\boldsymbol{J}_{i}$ into a centre-of-mass velocity field $\boldsymbol{v}=\sum_{i=0}^{N_{\mathrm{c}}}\boldsymbol{J}_{i}$ and relative fluxes $\boldsymbol{j}_{i}=\phi_{i}(\boldsymbol{v}_{i}-\boldsymbol{v})$, where $\boldsymbol{v}_{i}=\boldsymbol{J}_{i}/\phi_{i}$ is the velocity of component $i$ in the lab frame. Since many biomolecules have mass densities close to water, we here assumed equal mass density for all components, and refer the reader to ref. [101] for the general case. Hence,
@@
$\partial_{t}\boldsymbol{v}+\rho\boldsymbol{v}\!\cdot\!\nabla\boldsymbol{v}=\nabla\cdot\boldsymbol{\sigma}\;,$ (2.9)
@@
The coupled partial differential equations (2.14) for the fields $\{\phi_{i}\}$ typically permit stationary homogeneous solutions. For instance, without chemical reactions ($s_{i}=0$), each homogeneous state is stationary because of material conservation. The stability of such states can be assessed using linear stability analysis, which involves linearizing equations (2.14) and (2.7). In particular, the homogeneous state $\phi_{i}(\boldsymbol{r})=\bar{\phi}_{i}$ is stable if the associated Jacobian
$\mathcal{J}_{ij}$ $=-\nu_{0}\Omega_{i}^{2}\boldsymbol{q}^{2}\sum_{l=1}^{N_{\mathrm{c}}}\Lambda_{il}\left[\frac{\partial^{2}f}{\partial\phi_{l}\phi_{j}}+\kappa_{lj}\boldsymbol{q}^{2}\right]+\frac{\partial s_{i}}{\partial\phi_{j}}$ (2.15)
has only negative eigenvalues for all wave vectors $\boldsymbol{q}$. In contrast, the homogeneous state is unstable if $\mathcal{J}_{ij}$ has positive eigenvalues and the largest eigenvalue will typically determine the dominant length scale $2\pi/|\boldsymbol{q}|$ from the associated wavevector $\boldsymbol{q}$. However, since equations (2.14) are nonlinear, such a stability analysis only provides information about the initial dynamics.
@@ 2.1.3 Behaviour of binary liquids
Figure 2.1: Phase behaviour of a binary liquid. (A) Free energy density $f$ given by equation (2.16) as a function of the fraction $\phi$ (blue line) for $\Omega=1$, $\chi=2.5$, and $w=0.2$. The common-tangent construction (black line) solves equations (2.30) and thus determines the equilibrium fractions $\phi_{\mathrm{out}}$ and $\phi_{\mathrm{in}}$ (black disks) without Laplace pressure $P_{\gamma}$. Finite $P_{\gamma}$ require two parallel tangents offset by $P_{\gamma}$ (gray lines). The inflection points (orange disks) enclose the spinodal region where $f^{\prime\prime}(\phi)<0$. (B) Exchange chemical potential $\bar{\mu}$ given by equation (2.17) as a function of $\phi$ corresponding to panel A. The blue regions of equal area indicating a Maxwell construction, equivalent to the common-tangent construction. (C) Osmotic pressure $\Pi=\nu^{-1}\phi\bar{\mu}-f$ as a function of $\phi$ corresponding to panel A. (D) $\Pi$ as a function of $\bar{\mu}$ corresponding to panel A. (E, F) Phase diagrams corresponding to panel A highlighting the equilibrium volume fractions (black binodal line) and the spinodal region (within the orange spinodal line) for a symmetric mixture ($\Omega=1$, panel E) and for large solutes ($\Omega=10$, panel F). The critical point is marked by a star and grey areas denotes the region where bi-continuous structures have lower energy than droplets in two dimensions (obtained by comparing interfacial energies of disks and stripes for equal length scales). (G, H) Grand-canonical phase diagram corresponding to panels E and F. The binodals (black lines) mark the first-order transition between dilute and dense phases. Both phases are linearly stable in the spinodal region (between orange dotted lines).
@@
and determines the phase behaviour of the liquid; see figure 2.1. Note that the internal energy $w$ only shifts the chemical potentials and as such does not influence phase behavior; it will become important when discussing chemical reactions in section 5.
A homogeneous system with volume fraction $\phi$ is stable with respect to small perturbations when the Jacobian given in equation (2.15) is negative, i.e., if $f^{\prime\prime}(\phi)\propto\bar{\mu}^{\prime}(\phi)>0$}. This occurs when the interaction parameter $\chi$ is smaller than
@@
[…]
@@
for $\Omega=1$. In the limit of strong interactions ($\chi\gg 1$), this predicts $\phi^{(0)}_{\mathrm{out}}\approx e^{-\chi}$, which reveals that even a minute increase of the interaction energy by $2k_{\mathrm{B}}T$ (so $\chi$ is increased by $2$) leads to an almost $10$-fold decrease in the saturation concentration $\phi^{(0)}_{\mathrm{out}}$. In contrast, the spinodal branch scales as $\phi^{\mathrm{spin}}_{\mathrm{out}}\approx 1/(2\Omega\chi)$, indicating that a linear stability analysis of the homogeneous state is inadequate to study phase separation in detail.
[…]
@@
which implies an extra term $\kappa\nabla^{2}\phi$ in the exchange chemical potential $\bar{\mu}$ given in equation (2.17). The kinetics of the binary liquid follow from equation (2.14),